Two electrons are separated by 1.00 nm .
(a) What is the magnitude of the electric force each electron exerts on the other?
(b) Is this force attractive or repulsive?
1
Expert's answer
2021-08-24T06:52:21-0400

(a) The magnitude of the electric force each electron exerts on the other is equal

"F=k\frac{e^2}{r^2}\\
=9*10^9\:\rm N\cdot m^2/C^2*\frac{(1.6*10^{-19}\:\rm C)^2}{(1.0*10^{-9}\:\rm m)^2}=2.3*10^{-10}\:\rm N"

(b) Repulsive